10CM TO THE 3RD POWER OF A NORMAL BLOOD SPECIMEN CONTAINS 1.2G OF HEMEGLOBIN. HOW MANY GRAMS DOES 18CM TO THE 3RD POWER OF THE SAME BLOOD CONTAIN?
aLSO CAN YOU TELL ME HOW TO WRITE THE EQUATIONS LIKE 3 TO THE 3RD POWER AS WELL AS OTHER EQUATIONS THANK YOU.

Answer by stanbon(75887)

10CM TO THE 3RD POWER OF A NORMAL BLOOD SPECIMEN CONTAINS 1.2G OF HEMEGLOBIN. HOW MANY GRAMS DOES 18CM TO THE 3RD POWER OF THE SAME BLOOD CONTAIN?
------
Equation:
(10cm)^3 has 1.2G of Hemeglobin.
----
(18cm)^3 has xG of Hemeglobin.
----
Solve for "x":
Using a proportion:
x/(18^3) = 1.2/10^3
---
x = (18^3)(1.2/10^3)
---
x = 1.8^3*1.2
x = 5.832*1.2
x = 6.9984G
=============================
Cheers,
Stan H.
=============================